OpenStudy (anonymous):

Simplify the following expression. 3y-2x+z-2x+3(x+y) A. -x+4y+z B. -x+6y+z C. x+4y+z D. 3x+4y+z

OpenStudy (anonymous):

i think its c.

OpenStudy (kropot72):

First expand 3(x + y) to give 3(x + y) = 3x + 3y Then group the terms in x and the terms in y to give 3y + 3y + 3x - 2x -2x + z Can you find the answer now by adding the terms in x and adding the terms in y?

OpenStudy (anonymous):

no so it will be B.

OpenStudy (anonymous):

6y-x+z i feel like i got that wrong

OpenStudy (anonymous):

NO CUZ LIKE THAT GURL SAID 3Y+3Y= WOULD EQUAL 6Y

OpenStudy (kropot72):

@brittbrat213 You got it right. When you rearrange the terms in your answer you get B.

OpenStudy (anonymous):

oh okay i get it now will it be the same when it uses exponents? or will it be different

OpenStudy (kropot72):

It would be best if you give a question with exponents. You must group terms with the same variable and the same exponent.

OpenStudy (anonymous):

i have trouble on this 2 sometimes dont feel bad

OpenStudy (anonymous):

Simplify the expression. 2x^2+8+3x^2+4x-10 A. 5x^2+4x-2 B. 2x^2+x-10 C. 2x^2-x+10 D. 5x^2+7x-11

OpenStudy (anonymous):

i think B. or D.

OpenStudy (kropot72):

\[2x ^{2}+3x ^{2}+4x+8-10=?x ^{2}+4x-2\] Can you find the coefficient of x^2 ?

OpenStudy (anonymous):

is it 5?

OpenStudy (kropot72):

Yes, the coefficient of x^2 is 5. So the correct choice is A :)

OpenStudy (anonymous):

can you help me with 3 more questions? :s

OpenStudy (kropot72):

Best if you post them as new questions so others can help too :)